A country has the following information:


Autonomous consumption is 500

Marginal propensity to save is 0.2

Autonomous investment is 800

Marginal propensity to invest equals 0.12

Autonomous tax is zero

The marginal tax rate is 0.15

Government spending is equal to 1000

Export is 400

Autonomous Import is equal to 200

Marginal propensity to import is equal to 0.05

Question:

a.                            Calculate equilibrium output, trade balance, and budget balance. Calculate the total savings of the economy.

b.                           Calculate the amount of government expenditures or taxes that need to be adjusted knowing the potential output is 9000

c.                            If the marginal tax rate is the same, then what is value of government spending in order to balance the government’s budget?

d.                          If the government wants to spend 200 more but do not want to make output change, how much taxes should be changed?

e.                            If the government increases the spending 50 and increase the tax 50, how much the aggregate demand and output change?


1
Expert's answer
2021-03-14T19:06:14-0400

a. The equilibrium output is:

"Y = C + I + G + NX = 500 + 0.8\u00d7(1 - 0.15)Y + 800 + 0.12Y + 1000 + 400 - (200 + 0.05Y) = 0.75Y + 2500,"

0.25Y = 2500,

Y = 10000.

Trade balance is:

400 - (200 + 0.05×10000) = -300.

Budget balance is:

0.15×10000 - 1000 = 500.

Total savings of the economy are:

S = 10000 - (500 + 0.8×(1 - 0.15)×10000) - 1000 = 1700.

b. If the potential output is 9000, then either government expenditures should be decreased or taxes increased to close the inflationary gap.

c. If the marginal tax rate is the same, then the value of government spending should be increased by 500 in order to balance the government’s budget.

d. If the government wants to spend 200 more but do not want to make output change, then the tax rate should be increased by 0.02.

e. If the government increases the spending by 50 and increase the tax by 50, then the aggregate demand and output change will increase.